# Generating quantum non-local entanglement with mechanical rotations

### v1

@@ Abstract
Recent experiments have searched for evidence of the impact of non-inertial motion on the entanglement of particles. The success of these endeavours has been hindered by the fact that such tests were performed within spatial scales that were only "local" when compared to the spatial scales over which the non-inertial motion was taking place. We propose a Sagnac-like interferometer that, by challenging such bottlenecks, is able to achieve entangled states through a mechanism induced by the mechanical rotation of a photonic interferometer. The resulting states violate the Bell-Clauser-Horne-Shimony-Holt (CHSH) inequality all the way up to the Tsirelson bound, thus signaling strong quantum nonlocality. Our results demonstrate that mechanical rotation can be thought of as resource for controlling quantum non-locality with implications also for recent proposals for experiments that can probe the quantum nature of curved spacetimes and non-inertial motion.
@@ I Introduction
More recently, photonic technologies have enabled the exploration of rotation-induced quantum phenomena with two-photon experiments. Polarization-entangled photon pairs were shown to be robust against a 30g acceleration achieved on a rotating centrifuge [23]. Using a Hong-Ou-Mandel interferometer on a rotating platform it was found that low frequency mechanical rotations affect bunching statistics [24]. Super-resolution and Sagnac phase sensitivity beyond the shot-noise limit was achieved in [25] using path-entangled NOON states, and milli-radian phase resolution was achieved in [26], allowing the measurement of the Earth’s angular frequency of $\sim 10\mu\text{Hz}$. Furthermore, it was suggested that photonic entanglement can be revealed or concealed using non-inertial motion accessible to current experiments [27]. Using a Hong-Ou-Mandel interferometer with nested arms it was demonstrated that photonic behavior can change from bunching to antibunching (i.e., from bosonic to fermionic) solely due to mechanical rotations [28]. Moreover, it was shown that rotational motion can change the phase of polarization entangled states enabling transitioning between pairs of Bell states [29].
A step further is to demonstrate the actual generation of entanglement using non-inertial motion. The approach proposed in Ref. [30] made use of a multi-path Sagnac interferometer to achieve a maximally entangled path-polarization state of a single-photon. Such state would be suitable for quantum non-contextuality tests aimed at ascertaining whether observables can be assigned preexisting values prior to measurements [31]. In principle, the generated entanglement could be also transferred to two spatially separated physical systems [32], but the experimental demonstration of such procedures remains experimentally challenging [33, 34]. It is thus not immediately obvious whether the single-photon scheme proposed in  [30] would allow to unambiguously demonstrate the generation of genuine quantum (nonlocal) entanglement as opposed to local entanglement.
@@
Inspired by these two-party schemes, in this work we show that it is possible to use mechanical rotations to generate and control nonlocal quantum entanglement in an experimental regime that is fully accessible to current photonic technology. We propose a simple scheme consisting of a single Sagnac fibre loop, linear optics elements, photon-pair sources, and a Bell-test detection setup, all placed on a rotating platform. We find that an initially separable two-photon state can be transformed into a maximally entangled state of polarization that violates significantly the Bell-Clauser-Horne-Shimony-Holt (CHSH) inequality [40] as a function of the angular frequency of rotation, obtaining a simple formula for the frequency required to saturate Tsirelson’s bound [41].
Figure 1: (a) Photonic setting for the rotation-controlled generation of quantum non-local states of polarization. Two photons are initially prepared in the separable state $|HV\rangle$ and injected into the setup. Beam-splitters BS1 and BS2 send the photons into a Sagnac loop, and then redirect them towards two individual detection stages. As the photons entering the Sagnac loop through Beam-splitter BS3 are initially prepared in orthogonal polarization states, they do not interact at any point via electromagnetic couplings. We measure the polarization of the photons with a standard Bell detection scheme. One photon is measured using the setup at the top (managed by Alice) and one photon is measured with the setup on the right (managed by Bob). (b) Theoretical prediction of the violation of the Bell-CHSH inequality (blue) and the probability of detection (orange) [cf. Eq. (11) and (13), respectively]. For concreteness, we have set the photon wavelength to $\lambda=1\,\text{$\mu$m}$ ($\omega=2\pi c/\lambda$ ) and the interferometric area to $\sim 7.8\,\text{m}^{2}$ (e.g., $10$ loops of fiber with radius $r=0.5\,\text{m}$ with a total length of $\sim 31.5\,\text{m}$). The loss of a photon pair occurs when the beam-splitters do not direct the photons to Alice and Bob, thus lowering the probability of detection. We predict that a violation ($|{S}|>2$) occurs periodically with the rotation frequency as stated by Eq. (12). The Tsirelson’s bound is first achieved for $\Omega_{\text{Bell}}\sim 0.4\,\text{Hz}$.
@@
We consider the setup shown in Fig. 1 (a). It consists of two separate lasers, each pumping a non-linear crystal that, through Type I spontaneous parametric down-conversion, generates photon pairs of the same polarization of which we keep just one photon. The signals thus generated are first directed by beam-splitters (BSs) into a Sagnac loop, and then towards two spatially separated Bell-detection apparatuses. We assume that two independent photons are initially prepared in the separable state:
@@
where $\hat{a}_{H}$ ($\hat{b}_{V}$) denotes horizontal H (vertical V) polarization mode. We assume that the two photons propagate past beamsplitters BS1 and BS2 retaining the form of the state in Eq. (1). As the two photons enter the Sagnac loop through BS3, the state changes according to a beam-splitter transformation into
@@
|$\ket{\psi_{1}}{\to}\ket{\psi_{2}}{=}\frac{1}{2}\left(e^{i\frac{\phi}{2}}\hat{a}_{H}^{\dagger}{+}ie^{-i\frac{\phi}{2}}\hat{b}_{H}^{\dagger}\right)\left(ie^{i\frac{\phi}{2}}\hat{a}_{V}^{\dagger}{+}e^{-i\frac{\phi}{2}}\hat{b}_{V}^{\dagger}\right)\!|0\rangle,$ (3)
@@
As the photons exit the central loop through BS3, we apply the inverse beam-splitter transformation to Eq. (3), giving us
$\frac{1}{4}$ $\left[\left(e^{i\phi/2}(-i\hat{a}_{H}^{\dagger}+\hat{b}_{H}^{\dagger})+ie^{-i\phi/2}(\hat{a}_{H}^{\dagger}-i\hat{b}_{H}^{\dagger})\right)\right.$ (5)
×$\times\left.\left(ie^{i\phi/2}(-i\hat{a}_{V}^{\dagger}+\hat{b}_{V}^{\dagger})+e^{-i\phi/2}(\hat{a}_{V}^{\dagger}-i\hat{b}_{V}^{\dagger}\right)\right]|0\rangle,$
which, after rearranging, gives
$\left(\frac{1}{2}\left[(\text{cos}\phi-1)\hat{a}_{H}^{\dagger}\hat{b}_{V}^{\dagger}+(\text{cos}\phi+1)\hat{b}_{H}^{\dagger}\hat{a}_{V}^{\dagger}\right]\right.$
$\left.+\frac{\text{sin}\phi}{2}\left[\hat{a}_{H}^{\dagger}\hat{a}_{H}^{\dagger}-\hat{b}_{V}^{\dagger}\hat{b}_{V}^{\dagger}\right]\right)$ $|0\rangle.$ (6)
The first line of Eq. (6) represents the case when each pair of detectors detects one photon, i.e., Alice detects the mode $a$ and Bob detects mode $b$ or vice versa, while the second line represents the case when both photons arrive at the same pair of detectors, i.e., either both to Alice or both to Bob. As we want to compute quantum correlations between Alice and Bob we consider the state conditional on the postselection of the events that provide coincidences at the detectors, which reads
$|\psi_{\text{f}}\rangle$ $=\frac{\text{cos}(\phi)+1}{\sqrt{2(\text{cos}(\phi)^{2}+1)}}|H\,V\rangle+\frac{\text{cos}(\phi)-1}{\sqrt{2(\text{cos}(\phi)^{2}+1)}}|V\,H\rangle,$ (7)
where we have used $\hat{a}_{H}^{\dagger}\hat{b}_{V}^{\dagger}|0\rangle\equiv|H\,V\rangle$ and $\hat{a}_{V}^{\dagger}\hat{b}_{H}^{\dagger}|0\rangle\equiv|V\,H\rangle$, and included the overall state normalization for completeness.
$|\psi_{\text{f}}\rangle=\frac{1}{\sqrt{2}}\left(|H\,V\rangle-|V\,H\rangle\right),$ (8)
which is usually denoted as the $|\Psi^{-}\rangle$ state.
Our results depend critically on two steps: (i) on the post-selection step from Eq. (6) to (7), where we have discarded the photons that are not going to Alice and Bob, and (ii) on a non-zero rotationally induced phase $\phi\neq 0$, which arises only for non-zero frequencies of rotation $\Omega\neq 0$. Importantly, the post-selection step is not enough to induce entanglement in the absence of mechanical rotation as discussed above for the case $\phi=0$.
@@ III Violation of the Bell-chsh inequality
where
$E(\bm{a},\bm{b})=\langle\psi_{\text{f}}|(\bm{a}\cdot\bm{\sigma})\otimes(\bm{b}\cdot\bm{\sigma})|\psi_{\text{f}}\rangle$ (10)
denote quantum correlation functions. Here, $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ is the vector of the Pauli matrices and $\bm{a}$,$\bm{a}^{\prime}$,$\bm{b}$,$\bm{b}^{\prime}$ are vectors whose direction determines the value of $S$. In order to achieve the highest violation of the Bell test, we choose the vectors $\bm{a}=(1,0,0)$, $\bm{a}^{\prime}=(0,1,0)$, $\bm{b}=(1,1,0)/\sqrt{2}$, $\bm{b}^{\prime}=(-1,1,0)/\sqrt{2}$, thus obtaining

### v2

@@ Abstract
Recent experiments have searched for evidence of the impact of non-inertial motion on the entanglement of particles. The success of these endeavours has been hindered by the fact that such tests were performed within spatial scales that were only "local" when compared to the spatial scales over which the non-inertial motion was taking place. We propose a Sagnac-like interferometer that, by challenging such bottlenecks, is able to achieve entangled states through a mechanism induced by the mechanical rotation of a photonic interferometer. The resulting states violate the Bell-Clauser-Horne-Shimony-Holt (Bell-CHSH) inequality all the way up to the Tsirelson bound, thus signaling strong quantum nonlocality. Furthermore, we show that the Bell-CHSH inequality remains violated even without using any form of post-selection up to the value $1+\sqrt{2}$. Our results demonstrate that mechanical rotation can be thought of as resource for controlling quantum non-locality with implications also for recent proposals for experiments that can probe the quantum nature of curved spacetimes and non-inertial motion.
@@ I Introduction
More recently, photonic technologies have enabled the exploration of rotation-induced quantum phenomena with two-photon experiments. Polarization-entangled photon pairs were shown to be robust against a 30g acceleration achieved on a rotating centrifuge [23]. Using a Hong-Ou-Mandel interferometer on a rotating platform it was found that low frequency mechanical rotations affect bunching statistics [24]. Super-resolution and Sagnac phase sensitivity beyond the shot-noise limit was achieved in [25] using path-entangled NOON states, and milli-radian phase resolution was achieved in [26], allowing the measurement of the Earth’s angular frequency of $\sim 10\mu\text{Hz}$. Furthermore, it was suggested that photonic entanglement can be revealed or concealed using non-inertial motion accessible to current experiments [27]. Using a Hong-Ou-Mandel interferometer with nested arms it was demonstrated that photonic behavior can change from bunching to antibunching (i.e., from bosonic to fermionic) solely due to mechanical rotations [28]. Moreover, it was shown that rotational motion can change the phase of polarization entangled states enabling transitioning between pairs of Bell states [29]. Crucially, these experiments involve quantum states that are entangled prior to rotation and rotation is only used to modify, probe, or enhance existing quantum features.
A significant conceptual step further is to demonstrate the actual generation of entanglement using non-inertial motion. The approach proposed in Ref. [30] made use of a multi-path Sagnac interferometer to achieve a maximally entangled path-polarization state of a single-photon. Such state would be suitable for quantum non-contextuality tests aimed at ascertaining whether observables can be assigned preexisting values prior to measurements [31]. In principle, the generated entanglement could be also transferred to two spatially separated physical systems [32], but the experimental demonstration of such procedures remains experimentally challenging [33, 34]. It is thus not immediately obvious whether the single-photon scheme proposed in  [30] would allow to unambiguously demonstrate the generation of genuine quantum (nonlocal) entanglement as opposed to local entanglement.
@@
Inspired by these two-party schemes, in this work, we propose an experimentally viable scheme - fully accessible to current photonic technology - in which rotation itself is the mechanism that generates nonlocal entanglement between two initially unentangled photons. This marks a shift, from manipulating the entanglement to demonstrating its rotationally induced creation. Our scheme consists of a single Sagnac fibre loop, linear optics elements, photon-pair sources, and a Bell-test detection setup, all placed on a rotating platform. We find that an initially separable two-photon state can be transformed into a maximally entangled state of polarization that violates significantly the Bell-Clauser-Horne-Shimony-Holt (Bell-CHSH) inequality [40] as a function of the angular frequency of rotation, obtaining a simple formula for the frequency required to saturate Tsirelson’s bound [41]. Furthermore, we show that even without any form of post-selection, the Bell-CHSH inequality remains violated for the same angular frequencies of rotation, achieving the maximum value $1+\sqrt{2}$ [42]. We conclude by briefly discussing the interpretation of the uncovered link between mechanical motion and quantum nonlocality.
Figure 1: (a) Photonic setting for the rotation-controlled generation of quantum non-local states of polarization. Two photons are initially prepared in the separable state $|HV\rangle$ and injected into the setup. The circulators, denoted by C, first send the photons into a Sagnac loop, and then redirect them towards two individual detection stages (only the directional paths, $1$ to $2$ and $2$ to $3$, are allowed by the circulators). As the photons entering the Sagnac loop through Beam-splitter BS are initially prepared in orthogonal polarization states, they do not interact at any point via electromagnetic couplings. We measure the polarization of the photons with a standard Bell detection scheme. (b) Theoretical prediction of the violation of the Bell-CHSH inequality. With Post-selection refers to the case where one photon is measured using the setup at the top (managed by Alice) and one photon is measured with the setup on the right (managed by Bob). Without Post-selection refers to the case where all the photons are considered in the analysis, which includes the detection of one photon by Alice and one of photon by Bob as well as of both photons by either Alice or Bob. For concreteness, we have set the photon wavelength to $\lambda=1\,\text{$\mu$m}$ ($\omega=2\pi c/\lambda$) and the interferometric area to $\sim 7.8\,\text{m}^{2}$ (e.g., $10$ loops of fiber with radius $r=0.5\,\text{m}$ with a total length of $\sim 31.5\,\text{m}$). We predict that a violation ($|{S}|>2$) occurs periodically with the rotation frequency as stated by Eq. (13). The maximum violation is first achieved for $\Omega_{\text{Bell}}\sim 0.4\,\text{Hz}$, and the condition $|S|>2$ is achieved in the interval $\Omega\in 0.4\pm 0.1$ Hz (see Appendix A for more details).
@@
We consider the configuration shown in Fig. 1 (a), where two single photons are propagating through the system. The full setup involves a pulsed laser source (not shown), pumping two spatially separated type-I non-linear crystals, each producing photon pairs via spontaneous parametric down-conversion (SPDC). One photon from each pair is detected to herald the presence of its twin, resulting in two heralded single photons that are mutually uncorrelated, as they originate from independent SPDC processes. These signals are first directed into a Sagnac loop through circulators and a beam-splitter (BS), and then towards two spatially separated Bell-detection apparatuses. We assume that two independent photons are initially prepared in the separable state:
@@
where $\hat{a}_{H}$ ($\hat{b}_{V}$) denotes horizontal H (vertical V) polarization mode.
As the two photons enter the Sagnac loop through BS, the state changes according to a beam-splitter transformation into
@@
|$\ket{\psi_{1}}{\to}\ket{\psi_{2}}{=}\frac{1}{2}\left[e^{i\frac{\phi}{2}}\hat{a}_{H}^{\dagger}{+}ie^{-i\frac{\phi}{2}}\hat{b}_{H}^{\dagger}\right]\left[ie^{i\frac{\phi}{2}}\hat{a}_{V}^{\dagger}{+}e^{-i\frac{\phi}{2}}\hat{b}_{V}^{\dagger}\right]\!|0\rangle,$ (3)
@@
As the photons exit the central loop through the BS, we apply the inverse beam-splitter transformation to Eq. (3), giving us
|$\ket{\psi_{2}}$ |${\to}\ket{\psi_{\text{tot}}}{=}$ (5)
$=\frac{1}{4}\left[\left(e^{i\phi/2}(-i\hat{a}_{H}^{\dagger}+\hat{b}_{H}^{\dagger})+ie^{-i\phi/2}(\hat{a}_{H}^{\dagger}-i\hat{b}_{H}^{\dagger})\right)\right.$
×$\times\left.\left(ie^{i\phi/2}(-i\hat{a}_{V}^{\dagger}+\hat{b}_{V}^{\dagger})+e^{-i\phi/2}(\hat{a}_{V}^{\dagger}-i\hat{b}_{V}^{\dagger}\right)\right]|0\rangle.$
After rearranging the terms in Eq. (5) we find
$|\psi_{\text{tot}}\rangle=$ $\frac{1}{2}\sqrt{3+\text{cos}(2\phi)}|\psi_{\text{\text{\text{fav}}}}\rangle+\frac{\text{sin}\phi}{\sqrt{2}}|\psi_{\text{\text{\text{unfav}}}}\rangle,$ (6)
where we have introduced the normalized states
$|\psi_{\text{\text{\text{fav}}}}\rangle=\left[\frac{(\text{cos}\phi-1)\hat{a}_{H}^{\dagger}\hat{b}_{V}^{\dagger}}{\sqrt{3+\text{cos}(2\phi)}}+\frac{(\text{cos}\phi+1)\hat{b}_{H}^{\dagger}\hat{a}_{V}^{\dagger}}{\sqrt{3+\text{cos}(2\phi)}}\right]|0\rangle,$ (7)
$|\psi_{\text{\text{unfav}}}\rangle=\frac{1}{\sqrt{2}}\left[\hat{a}_{H}^{\dagger}\hat{a}_{V}^{\dagger}-\hat{b}_{H}^{\dagger}\hat{b}_{V}^{\dagger}\right]|0\rangle.$ (8)
$|\psi_{\text{\text{\text{fav}}}}\rangle$ represents the case when each pair of detectors detects one photon, i.e., Alice detects the mode $a$ and Bob detects mode $b$ or vice versa, while $|\psi_{\text{\text{\text{unfav}}}}\rangle$ represents the case when both photons arrive at the same pair of detectors, i.e., either both to Alice or both to Bob. As we want to compute quantum correlations between Alice and Bob we will refer to $|\psi_{\text{\text{\text{fav}}}}\rangle$ and $|\psi_{\text{\text{\text{unfav}}}}\rangle$ as the favorable and unfavorable state, respectively.
@@ III Violation of the Bell-CHSH inequality
where the quantum correlations are given by
$E(\bm{v},\bm{u})=\langle\psi|\hat{A}_{\bm{v}}\hat{B}_{\bm{u}}|\psi\rangle,$ (10)
$|\psi\rangle$ is the state of the system, and $\hat{A}_{\bm{v}}$ and $\hat{B}_{\bm{u}}$ correspond to the measurements performed by Alice and Bob, respectively ($\bm{v}=\bm{a}$,$\bm{a}^{\prime}$ and $\bm{u}=\bm{b}$,$\bm{b}^{\prime}$ denote unit vectors parametrizing the measurement operators). The four terms appearing in Eq. (9) are assocated to four different joint measurements performed by Alice and Bob. Local realism enforces the Bell-CHSH inequality $|S|\leq 2$. However, as we will see, suitable choices of the angular frequency of rotation, encoded in the phase $\phi$ in Eq. (6), allows to violate such constraint. We show below that the Bell-CHSH inequality is violated up to the value $|S|=1+\sqrt{2}$ by considering the total state in Eq. (6), establishing the phenomena of rotationally induced quantum non-locality (Sec. III.1). Furthermore, we show that Tsirelson’s bound $|S|=2\sqrt{2}$ can be achieved by post-selecting coincidences at the detectors (Sec. III.2).
III.1 Non-locality without post-selection
We define the observables $\hat{A}_{\bm{u}}$ and $\hat{B}_{\bm{v}}$ by prescribing how they act on the basis states of our Hilbert space (we recall that observables, which are linear maps, can always be defined in this way). The total Hilbert space, and in particular its basis, can be read out from the total final state $|\psi_{\text{tot}}\rangle$ in Eqs. (6)-(8). On the Hilbert subspace spanned by $\hat{a}_{H}^{\dagger}\hat{b}_{V}^{\dagger}|0\rangle$ and $\hat{b}_{H}^{\dagger}\hat{a}_{V}^{\dagger}|0\rangle$ (i.e., the basis states of the favourable subspace) we define $\hat{A}_{\bm{a}}\equiv\bm{a}\cdot\bm{\sigma}$ and $\hat{B}_{\bm{b}}\equiv\bm{b}\cdot\bm{\sigma}$, where $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ is the vector of the Pauli matrices. In order to achieve the highest violation of the Bell test, we choose the vectors $\bm{a}=(1,0,0)$, $\bm{a}^{\prime}=(0,1,0)$, $\bm{b}=(1,1,0)/\sqrt{2}$, $\bm{b}^{\prime}=(-1,1,0)/\sqrt{2}$. On the Hilbert subspace spanned by $\hat{a}_{H}^{\dagger}\hat{a}_{V}^{\dagger}|0\rangle$ and $\hat{b}_{H}^{\dagger}\hat{b}_{V}^{\dagger}|0\rangle$ (i.e., the basis states of the unfavourable subspace) we define $\hat{A}_{\bm{a}}=\mathbb{I}$ and $\hat{B}_{\bm{a}}=\mathbb{I}$, where $\mathbb{I}$ denotes the identity operator [42].
To compute the Bell-CHSH value we consider the total state without any post-selection, i.e., we set $|\psi\rangle\equiv|\psi_{\text{tot}}\rangle$. Inserting Eq. (6) into Eqs. (9) we eventually find:
$S=\left(1+\sqrt{2}\right)\sin^{2}(\phi).$ (11)
In particular, by setting $\phi=\pi/2+k\pi\leavevmode\nobreak\ (k\in\mathcal{Z})$, the state in Eq. (6) reduces to $|\psi_{\text{tot}}\rangle=\frac{1}{\sqrt{2}}(\psi_{\text{\text{\text{fav}}}}\rangle+|\psi_{\text{\text{\text{unfav}}}}\rangle)$, i.e.,
$|\psi_{\text{tot}}\rangle=\frac{1}{2}\left[-\hat{a}_{H}^{\dagger}\hat{b}_{V}^{\dagger}+\hat{b}_{H}^{\dagger}\hat{a}_{V}^{\dagger}+\hat{a}_{H}^{\dagger}\hat{a}_{V}^{\dagger}-\hat{b}_{H}^{\dagger}\hat{b}_{V}^{\dagger}\right]|0\rangle,$ (12)
and we achieve the maximum violation $|S|=1+\sqrt{2}$. This occurs when the mechanical frequency $\Omega$ takes the values
Here, $k<0$ ($k>0$) would correspond to an (anti-) clockwise sense of rotation.
The experimentally measurable Bell-CHSH correlation will be obtained by repeating the experiment with a large number of initial photon pairs in order to gather enough statistics. In Fig. 1(b), we plot the Bell-CHSH function for a set of values of the relevant physical parameters that are well within reach of existing photonic technology with previous experiments already achieving the required sensitivities [23, 27, 28, 25, 26, 29, 24].
Importantly, the Bell-CHSH violation depends critically only on the mechanical rotation. We recall that two initial state in Eq. (1) consists of photons that are in orthogonal polarization modes, and hence do not interact at the beam-splitter or via any electromagnetic interaction, and that no post-selection has been performed on the total final state in Eq. (6). The results captured by Eq. (9) thus depend critically only on the angular frequency of rotation $\Omega$: by tuning its value we can generate quantum non-locality, while without mechanical rotation no quantum non-locality can be established.
III.2 Tsirelson’s bound with postselection
We now consider the state conditional on the postselection of the events that provide coincidences at the detectors, i.e., quantum correlations between Alice and Bob, where we disregard the events arising from the unfavorable state $|\psi_{\text{unfav}}\rangle$ in Eq. (8). We discuss the probability of coincidence detection in Appendix B, which remains greater than $50\%$, and hence does not pose a severe limitation. We can rewrite the remaining favorable state $|\psi_{\text{fav}}\rangle$ defined in Eq. (7) as:
$|\psi_{\text{fav}}\rangle$ $=\frac{\text{cos}(\phi)+1}{\sqrt{3+\text{cos}(2\phi)}}|H\,V\rangle+\frac{\text{cos}(\phi)-1}{\sqrt{3+\text{cos}(2\phi)}}|V\,H\rangle,$ (14)
where we have introduced the commonly used notation $\hat{a}_{H}^{\dagger}\hat{b}_{V}^{\dagger}|0\rangle\equiv|H\,V\rangle$ and $\hat{a}_{V}^{\dagger}\hat{b}_{H}^{\dagger}|0\rangle\equiv|V\,H\rangle$.
$|\psi_{\text{fav}}\rangle=\frac{1}{\sqrt{2}}\left(|H\,V\rangle-|V\,H\rangle\right),$ (15)
which is usually denoted as the $|\Psi^{-}\rangle$ state.
To compute the Bell-CHSH value we set $|\psi\rangle\equiv|\psi_{\text{f}}\rangle$ in Eq. (9) to obtain the value